OpenStudy (anonymous):

Updated Batting Averages are computed using the formula: A = quantity of B plus K all over G A = batting average B = times at bat multiplied by player’s seasonal average K = hits in the new game G = times at bat plus times at bat in a new game Solve A = quantity of B plus K all over G for K. Please show your work. (6 points) Use your new formula and the values given in the table below to find the player’s hits in a new game (K). Round your answer to the nearest whole number. (4 points) Batter George 0.319 10.7 46 Jose 0.320 11.4 45 Dale 0.298 12.3 52 Michael 0.337 15.2 48

OpenStudy (anonymous):

George (A) 0.319 (B) 10.7 (G) 46 (K)? Jose (A) 0.320 (B) 11.4 (G) 45 ect....

OpenStudy (anonymous):

I'm pretty sure the answers are: George 4 Jose 3 Dale 3 Michael 1

OpenStudy (anonymous):

1) Solve A = quantity of B plus K all over G for K. A = ( B + K ) / G A * G = B + K K = A * G - B 2) Use your new formula and the values given in the table below to find the player’s hits in a new game (K). Let A = batting average Let B = times at bat multiplied by player’s seasonal average Let G = times at bat plus times at bat in a new game Let K = hits in the new game It is hard to make out the table, but I will presume the following: Player: George A = 0.319 B = 10.7 G = 46 Solving for K: K = A * G - B K = (0.319)(46) - 10.7 K = 14.674 - 10.7 K = 3.974 Rounded to the nearest whole number, K = 4 ----> George's hit in the a new game is 4. Player: Jose: A = 0.320 B = 11.4 G = 45 Solving for K: K = A * G - B K = (0.320)(45) - 11.4 K = 14.4 - 11.4 K = 3 ----> Jose's hits in a new game is 3. Player Dale: A = 0.298 Solving for K: K = A * G - B K = (0.298)(52) - 12.3 K = 15.496 - 12.3 K = 3.196 Rounded to the nearest whole number, K = 3 ----> Dale's hit in the a new game is 3. Player Michael A = 0.337 B = 15.2 G = 48 Solving for K: K = A * G - B K = (0.337)(48) - 15.2 K = 16.176 - 15.2 K = 0.976 Rounded to the nearest whole number, K = 1 ----> Michael's hit in the a new game is 1.

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

Yes, it helped me a ton! Thanks a lot! :)

OpenStudy (anonymous):

Did you get my message? @kathy0514

OpenStudy (anonymous):

no. wut message?

OpenStudy (anonymous):

I sent it again

OpenStudy (anonymous):

thanx